OpenStudy (anonymous):

MOTHER CHECKING SONS HW HE PUT 5 AND 2 IS THIS RIGHT EXPLAIN AND ANWER PLS What is the solution to the system of equations? x+y=5 x-2y=2

OpenStudy (littlebird):

x+y=5 ------(solving for y)-----> y=5-x x-2y=2 Substitution! x-2(5-x)=2 x-10+2x=2 3x=12 x=4 That boy is wrong!

OpenStudy (anonymous):

oh god he was so wrong so the anwer is 12 and 4? thank you im goin to have him correct it and try to explain those are the correct anwers ?